# theorem]Theorem theorem]Corollary theorem]Lemma theorem]Example theorem]Assumption theorem]Proposition

### v4

@@ 1 Introduction
Stochastic optimization methods such as SGD (Robbins & Monro, 1951b) or Adam (Kingma & Ba, 2014) are arguably at the core of the success of large-scale machine learning (LeCun et al., 2015; Schmidhuber, 2015). This success has led to significant (perhaps even excessive) research efforts dedicated to designing new variants of these methods (Schmidt et al., 2020). In all these methods, massive datasets are collected centrally on a server, and immense parallel computational resources of a data center are leveraged to perform training (Goyal et al., 2017; Brown et al., 2020). Meanwhile, modern machine learning is moving away from this centralized training setup with new paradigms emerging, such as i) distributed/federated learning, ii) semi-supervised learning, iii) personalized/multi-task learning, iv) model compression, Et cetera. Relatively little attention has been devoted to these more practical settings from the optimization community. In this work, we focus on extending the framework and tools of stochastic optimization to bear on these novel problems.
At the heart of these newly emergent training paradigms lies the following fundamental optimization question: We want to minimize a target loss function $f({\bm{x}})$, but computing its stochastic gradients is either very expensive or unreliable due to limited data. However, we assume having access to some auxiliary loss function $h({\bm{x}})$ whose stochastic gradient computation is relatively cheaper or more available. For example, in transfer learning, $f({\bm{x}})$ would represent the downstream task we care about and for which we have very little data available, whereas $h({\bm{x}})$ would be the pretraining task for which we have plenty of data (Yosinski et al., 2014). Similarly, in semi-supervised learning, $f({\bm{x}})$ would represent the loss over our clean labeled data, whereas $h({\bm{x}})$ represents the loss over unlabeled or noisily labeled data (Chapelle et al., 2009). Our challenge is the following question:
@@
Related work. Optimizing one function $f$ while accessing another function $h$ (or its gradients) is an important idea that has only been considered in specific cases in machine learning and optimization communities. To the best of our knowledge, this problem has never been regarded in all of its generality before this time. For this reason, we are limited to only citing works that used this idea in particular cases. Lately, masked training of neural networks was considered, for example, in (Alexandra et al., 2019; Amirkeivan et al., 2021); this approach is a special case of our framework, where the auxiliary information is given by the sub-network (or mask). In distributed optimization, (Shamir et al., 2013) define sub-problems based on available local information; the main problem with this approach is that the defined sub-problems need to be solved precisely in theory and to high precision in practice. In Federated Learning (Konecny et al., 2016; McMahan et al., 2017a; Mohri et al., 2019), the local functions (constructed using local datasets) can be seen as side information. Applying our framework recuperates an algorithm close to MiMe (Karimireddy et al., 2020a). In personalization, Chayti et al. (2021) study the collaborative personalization problem where one user optimizes its loss by using gradients from other available users (that are willing to collaborate); again, these collaborators can be seen as side information, one drawback of the approach in (Chayti et al., 2021) is that they need the same amount of work from the main function and the helpers, in our case we alleviate this by using the helpers more.
There is also auxiliary learning (Baifeng et al., ; Aviv et al., ; Xingyu et al., ) that is very similar to what we are proposing in this work. Auxiliary learning also has the goal of learning one given task using helper tasks, however, all these works come without any theoretical convergence guarantees, furthermore, our approach is more general.
The proposed framework is general enough to include all the above problems and more. More importantly, we don’t explicitly make assumptions on how the target function $f$ is related to the auxiliary side information $h$ (potentially a set of functions) like in Distributed optimization or Federated learning where we assume $f$ is the average of the side-information $h$. Also, it is not needed to solve the local problems precisely as expected by DANE (Shamir et al., 2013).
@@ 2 General Framework
Control variates and SVRG. (3) can be understood as a generalization of the control variate idea used in SVRG (Johnson & Zhang, 2013). To optimize a function $f({\bm{x}}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\bm{x}})$, SVRG uses the modified gradient ${\bm{g}}_{SVRG}=\nabla f_{i}({\bm{y}})-\nabla f_{i}({\bm{x}})+\nabla f({\bm{x}})$ where ${\bm{x}}$ is a snapshot that is updated less frequently and $i$ is sampled randomly so that this new gradient is still unbiased. The convergence of SVRG is guaranteed by the fact that the “error” of this new gradient is $\mathbb{E}\|{\bm{g}}_{SVRG}-\nabla f({\bm{x}})\|_{2}^{2}=\mathcal{O}(\|{\bm{y}}-{\bm{x}}\|_{2}^{2})$ so that if ${\bm{y}}-{\bm{x}}\rightarrow 0$, then convergence is guaranteed without needing to take small step-sizes. The main idea of our work is to use instead of $f_{i}$ another function $h$ that is related to $f$; this means using a gradient ${\bm{g}}=\nabla h({\bm{y}})-\nabla h({\bm{x}})+\nabla f({\bm{x}})$, then if we can still guarantee that the error is $\mathcal{O}(\|{\bm{y}}-{\bm{x}}\|_{2}^{2})$ everything should still work fine.
Other Variance reduction techniques. given the form of (3) that is very similar to the SVRG gradient, it is natural to ask what would happen when using a form similar to other variance reduction techniques such as SARAH (Nguyen et al., 2017) (this will amount to choosing a gradient ${\bm{g}}^{t}=\nabla h({\bm{y}}^{t})-\nabla h({\bm{y}}^{t-1})+{\bm{g}}^{t-1}$ with ${\bm{g}}^{0}=\nabla f({\bm{y}}^{0}:={\bm{x}})$). Unfortunately, this choice leads to the same theoretical rate obtained by the SVRG-like choice, the main reason being that on top of the biasedness of SARAH, the fact that $h$ is potentially different from $f$ (even in average) introduces another biasedness, which limits the potential gain; moreover, it is not evident how to treat the case where we only have access to stochastic gradients of $f$.
What if we can’t access the true gradient of $f$? When $f$ is not a finite average, we can’t access its true gradient; in this case, we propose replacing the “correction” $\nabla f({\bm{x}})-\nabla h({\bm{x}})$ by a quantity ${\bm{m}}_{f-h}$ that is a form of momentum (takes into account past observed gradients of $f-h$), the idea of using momentum is used to stabilize the estimate of the quantity $\nabla f({\bm{x}})-\nabla h({\bm{x}})$ as momentum can be used to reduce the variance. Specifically, we use ${\bm{g}}=\nabla h({\bm{y}})+{\bm{m}}_{f-h}$. We note that for this estimate we have $\mathbb{E}\|{\bm{g}}-\nabla f({\bm{y}})\|_{2}^{2}=\mathcal{O}(\|{\bm{y}}-{\bm{x}}\|_{2}^{2}+\|{\bm{m}}_{f-h}-\nabla f({\bm{x}})+\nabla h({\bm{x}})\|_{2}^{2})$ which means that ${\bm{g}}$ approximates $\nabla f({\bm{y}})$ as long as ${\bm{y}}$ is not far from ${\bm{x}}$ and ${\bm{m}}_{f-h}$ is a good estimate of the quantity $\nabla f({\bm{x}})-\nabla h({\bm{x}})$.
@@
General meta-algorithm. Based on the discussion above, we propose the (meta)-Algorithm 1: at the beginning of each round $t$, we have an estimate ${\bm{x}}^{t-1}$ of the minimizer. We sample a new $\zeta_{f-h}$, and compute ${\bm{g}}_{f-h}({\bm{x}}_{t-1},\zeta_{f-h})$, a noisy unbiased estimate of the gradient of $f-h$; we then update ${\bm{m}}_{f-h}^{t}$ a momentum of $f-h$. We transfer both ${\bm{x}}^{t-1}$ and ${\bm{m}}_{f-h}^{t}$ to the helper $h$ which uses both to construct a set of biased gradients ${\bm{d}}^{t}_{k}$ of $\nabla f({\bm{y}}^{t}_{k-1})$ that are updated by $h$. These biased gradients are then used to update the “local” states ${\bm{y}}^{t}_{k}$, then $f$ updates its own state ${\bm{x}}^{t}$ based on the last local states ${\bm{y}}^{t}_{0\leq k\leq K}$; throughout this work, we simply take ${\bm{x}}_{t}={\bm{y}}^{t}_{K}$.
0:  ${\bm{x}}_{0}$, $\eta$, $T$, $K$
@@
     update ${\bm{m}}_{f-h}^{t}\approx\nabla f({\bm{x}}^{t-1})-\nabla h({\bm{x}}^{t-1})$ momentum
@@
For our purposes we can take ${\bm{d}}^{t}_{k}={\bm{g}}_{h}({\bm{y}}^{t}_{k-1},\xi^{t,k}_{h})+{\bm{m}}^{t}_{f-h}$ which should be a good approximation of $\nabla f({\bm{y}}^{t}_{k-1})$ and ${\bm{m}}^{t}_{f-h}$ is a momentum of $f-h$, we will consider two options: classical momentum or MVR (for momentum based variance reduction (Cutkosky & Orabona, 2019)).
Decentralized auxiliary information. More generally, we can assume having access to $N$ auxiliary functions $h_{i}({\bm{x}}):=\mathbb{E}_{\xi_{h_{i}}}\big{[}h_{i}({\bm{x}};\xi_{h_{i}})\big{]}$. While we can treat this case by taking $h=(1/N)\sum_{i=1}^{N}h_{i}$, we propose a more interesting solution that also works if the helpers $h_{i}$ are decentralized and cannot live in the same server. In this case, we can sample a set $S^{t}$ of helpers, each $h_{i}\in S^{t}$ will do the updates exactly as in Algorithm1, but this time ${\bm{x}}^{t}$ will be constructed using all $\{{\bm{y}}^{t}_{i,K},i\in S^{t}\}$. In our case, we propose to use the average ${\bm{x}}^{t}=(1/S)\sum_{i\in S^{t}}{\bm{y}}^{t}_{i,K}$
@@ 3 Algorithms and Results
Naive approach. This approach is exactly as suggested by its name, naive; it simply ignores the part $f-h$ or, in other words, sets ${\bm{m}}^{t}_{f-h}=0$. The main idea is to use gradients (or gradient estimates) of $f$ whenever they are available and use gradients of $h$ when gradients of $f$ are not available. In our case, we alternate between one step using a gradient of $f$ and $(K-1)$-steps using gradients of $h$ without any correction. We will show that this approach suffers heavily from the bias between the gradients of $f$ and $h$. It is worth noting that in federated learning, this approach corresponds to Federated averaging (McMahan et al., 2017a).We remind the reader that we can take ${\bm{d}}^{t}_{k}={\bm{g}}_{h}({\bm{y}}^{t}_{k-1},\xi^{t,k}_{h})+{\bm{m}}^{t}_{f-h}$.
@@
where ${\bm{m}}_{f-h}^{t}$ is a momentum of $f-h$. We consider two methods for defining this momentum:
@@
AuxMOM simply uses the classical momentum, whereas AuxMVR uses the momentum-based variance reduction technique introduced in (Cutkosky & Orabona, 2019).
@@
In Assumption 3.2, we assume that we directly have access to unbiased gradient estimates of $f-h$; this does not restrict in any way our work since ${\bm{g}}_{f}-{\bm{g}}_{h}$ is such an estimate; however, this last estimate has a variance of $\sigma_{f-h}^{2}=\sigma_{f}^{2}+\sigma_{h}^{2}$, in general, it is possible to have a correlated estimate such that $\sigma_{f-h}^{2}<\sigma_{f}^{2}+\sigma_{h}^{2}$. If the batches $\xi_{f}$ and $\xi_{h}$ are drawn such that ${\bm{g}}_{f}$ and ${\bm{g}}_{h}$ are positively correlated, then it is even possible to have $\sigma_{f-h}^{2}<\sigma_{f}^{2}$.
@@
Discussion of the assumptions. Assumptions 3.1 and 3.2 are very common assumptions in the optimization literature. Under these two assumptions, it is well-known that SGD (Robbins & Monro, 1951a; Kiefer & Wolfowitz, 1952) has an optimal convergence rate. Assumption 3.3 was used extensively in Federated Learning (Karimireddy et al., 2020a; b) and was considered for personalization (Chayti et al., 2021).
Relaxing the Hessian similarity. Jingzhao et al. propose a generalized smoothness assumption where the norm of the Hessian can grow with the norm of the gradient. We believe it possible to extend our theory to accommodate such an assumption. In the same spirit, we can let $\|\nabla^{2}f({\bm{x}})-\nabla^{2}h({\bm{x}})\|_{2}$ grow with $\|\nabla f({\bm{x}})\|_{2}$. More important than this is finding similarity measures that apply to some of the potential applications we cite in Section 5.
@@ 4.1 Naive approach
In this section, we show the convergence results using the naive approach that uses a gradient of $f$ followed by $K-1$ gradients of $h$. For the analysis of this case (and only of this case), we need to make an assumption on the gradient bias between $f$ and $h$.
@@
There exists $f$ and $h$ satisfying assumptions 3.1,3.2,3.3,4.1 with $\delta=0$, $\sigma_{f}=\sigma_{h}=0$ such that $\frac{1}{KT}\sum_{t=1}^{T}\|\nabla f({\bm{x}}^{t-1})\|_{2}^{2}=\Omega(\zeta^{2})$. Using the biased SGD analysis in (Ajalloeian & Stich, 2020), it is easy to prove an upper bound, but we only need a lower bound for our purposes. Theorem 4.1 shows that this naive approach cannot guarantee convergence to less than $\zeta^{2}$ (up to some constant).
Note. Theorem 4.1 is very loose for Federated Averaging as, in this case, the function $f$ is directly tied to the helper entities (it is the average). Nevertheless, it’s worth noting that heterogeneity and client drift, which are akin to gradient bias, are recognized as factors that can constrain the performance of FedAVG. Consequently, there have been efforts to mitigate these issues through approaches such as those discussed in (Karimireddy et al., 2020b; a).

### v5

@@ 1 Introduction
Stochastic optimization methods such as SGD (robbins1951stochastic) or Adam (kingma2014adam) are arguably at the core of the success of large-scale machine learning (lecun2015deep; schmidhuber2015deep). This success has led to significant (perhaps even excessive) research efforts dedicated to designing new variants of these methods (schmidt2020descending). In all these methods, massive datasets are collected centrally on a server, and immense parallel computational resources of a data center are leveraged to perform training (goyal2017accurate; brown2020language). Meanwhile, modern machine learning is moving away from this centralized training setup with new paradigms emerging, such as i) distributed/federated learning, ii) semi-supervised learning, iii) personalized/multi-task learning, iv) model compression, Et cetera. Relatively little attention has been devoted to these more practical settings from the optimization community. In this work, we focus on extending the framework and tools of stochastic optimization to bear on these novel problems.
At the heart of these newly emergent training paradigms lies the following fundamental optimization question: We want to minimize a target loss function $f({\bm{x}})$, but computing its stochastic gradients is either very expensive or unreliable due to limited data. However, we assume having access to some auxiliary loss function $h({\bm{x}})$ whose stochastic gradient computation is relatively cheaper or more available. For example, in transfer learning, $f({\bm{x}})$ would represent the downstream task we care about and for which we have very little data available, whereas $h({\bm{x}})$ would be the pretraining task for which we have plenty of data (yosinski2014transferable). Similarly, in semi-supervised learning, $f({\bm{x}})$ would represent the loss over our clean labeled data, whereas $h({\bm{x}})$ represents the loss over unlabeled or noisily labeled data (chapelle2009semi). Our challenge is the following question:
@@
Related work. Optimizing one function $f$ while accessing another function $h$ (or its gradients) is an important idea that has only been considered in specific cases in machine learning and optimization communities. To the best of our knowledge, this problem has never been regarded in all of its generality before this time. For this reason, we are limited to only citing works that used this idea in particular cases. Lately, masked training of neural networks was considered, for example, in (ACDC; MaskedTraining); this approach is a special case of our framework, where the auxiliary information is given by the sub-network (or mask). In distributed optimization, (DANE) define sub-problems based on available local information; the main problem with this approach is that the defined sub-problems need to be solved precisely in theory and to high precision in practice. In Federated Learning (FL; MacMahan2017; Mohri2019), the local functions (constructed using local datasets) can be seen as side information. Applying our framework recuperates an algorithm close to MiMe (MiMe). In personalization, LinSpeedup study the collaborative personalization problem where one user optimizes its loss by using gradients from other available users (that are willing to collaborate); again, these collaborators can be seen as side information, one drawback of the approach in (LinSpeedup) is that they need the same amount of work from the main function and the helpers, in our case we alleviate this by using the helpers more.
There is also auxiliary learning (AuxTMDL; AuxLinImplicit; AdaptTaskReweighting) that is very similar to what we are proposing in this work. Auxiliary learning also has the goal of learning one given task using helper tasks, however, all these works come without any theoretical convergence guarantees, furthermore, our approach is more general.
The proposed framework is general enough to include all the above problems and more. More importantly, we don’t explicitly make assumptions on how the target function $f$ is related to the auxiliary side information $h$ (potentially a set of functions) like in Distributed optimization or Federated learning where we assume $f$ is the average of the side-information $h$. Also, it is not needed to solve the local problems precisely as expected by DANE (DANE).
@@ 2 General Framework
Control variates and SVRG. (3) can be understood as a generalization of the control variate idea used in SVRG (SVRG). To optimize a function $f({\bm{x}}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\bm{x}})$, SVRG uses the modified gradient ${\bm{g}}_{SVRG}=\nabla f_{i}({\bm{y}})-\nabla f_{i}({\bm{x}})+\nabla f({\bm{x}})$ where $\bm{x}$} is a snapshot that is updated less frequently and $i$ is sampled randomly so that this new gradient is still unbiased. The convergence of SVRG is guaranteed by the fact that the “error” of this new gradient is $\mathbb{E}\|{\bm{g}}_{SVRG}-\nabla f({\bm{x}})\|_{2}^{2}=\mathcal{O}(\|{\bm{y}}-{\bm{x}}\|_{2}^{2})$ so that if ${\bm{y}}-{\bm{x}}\rightarrow 0$, then convergence is guaranteed without needing to take small step-sizes. The main idea of our work is to use instead of $f_{i}$ another function $h$ that is related to $f$; this means using a gradient ${\bm{g}}=\nabla h({\bm{y}})-\nabla h({\bm{x}})+\nabla f({\bm{x}})$, then if we can still guarantee that the error is $\mathcal{O}(\|{\bm{y}}-{\bm{x}}\|_{2}^{2})$ everything should still work fine.
Other Variance reduction techniques. given the form of (3) that is very similar to the SVRG gradient, it is natural to ask what would happen when using a form similar to other variance reduction techniques such as SARAH (nguyen2017sarah) (this will amount to choosing a gradient ${\bm{g}}^{t}=\nabla h({\bm{y}}^{t})-\nabla h({\bm{y}}^{t-1})+{\bm{g}}^{t-1}$ with ${\bm{g}}^{0}=\nabla f({\bm{y}}^{0}:={\bm{x}})$). Unfortunately, this choice leads to the same theoretical rate obtained by the SVRG-like choice, the main reason being that on top of the biasedness of SARAH, the fact that $h$ is potentially different from $f$ (even in average) introduces another biasedness, which limits the potential gain; moreover, it is not evident how to treat the case where we only have access to stochastic gradients of $f$.
What if we can’t access the true gradient of $f$? When $f$ is not a finite average, we can’t access its true gradient; in this case, we propose replacing the “correction” $\nabla f({\bm{x}})-\nabla h({\bm{x}})$ by a quantity ${\bm{m}}_{f-h}$ that is a form of momentum (takes into account past observed gradients of fhf-h), the idea of using momentum is used to stabilize the estimate of the quantity $\nabla f({\bm{x}})-\nabla h({\bm{x}})$ as momentum can be used to reduce the variance. Specifically, we use ${\bm{g}}=\nabla h({\bm{y}})+{\bm{m}}_{f-h}$. We note that for this estimate we have $\mathbb{E}\|{\bm{g}}-\nabla f({\bm{y}})\|_{2}^{2}=\mathcal{O}(\|{\bm{y}}-{\bm{x}}\|_{2}^{2}+\|{\bm{m}}_{f-h}-\nabla f({\bm{x}})+\nabla h({\bm{x}})\|_{2}^{2})$ which means that $\bm{g}$} approximates $\nabla f({\bm{y}})$ as long as $\bm{y}$} is not far from $\bm{x}$} and ${\bm{m}}_{f-h}$ is a good estimate of the quantity $\nabla f({\bm{x}})-\nabla h({\bm{x}})$.
@@
General meta-algorithm. Based on the discussion above, we propose the (meta)-Algorithm 1: at the beginning of each round $t$, we have an estimate ${\bm{x}}^{t-1}$ of the minimizer. We sample a new $\zeta_{f-h}$, and compute ${\bm{g}}_{f-h}({\bm{x}}_{t-1},\zeta_{f-h})$, a noisy unbiased estimate of the gradient of fhf-h; we then update ${\bm{m}}_{f-h}^{t}$ a momentum of fhf-h. We transfer both ${\bm{x}}^{t-1}$ and ${\bm{m}}_{f-h}^{t}$ to the helper $h$ which uses both to construct a set of biased gradients ${\bm{d}}^{t}_{k}$ of $\nabla f({\bm{y}}^{t}_{k-1})$ that are updated by $h$. These biased gradients are then used to update the “local” states ${\bm{y}}^{t}_{k}$, then $f$ updates its own state ${\bm{x}}^{t}$ based on the last local states ${\bm{y}}^{t}_{0\leq k\leq K}$; throughout this work, we simply take ${\bm{x}}_{t}={\bm{y}}^{t}_{K}$.
[…]
@@
  update ${\bm{m}}_{f-h}^{t}\approx\nabla f({\bm{x}}^{t-1})-\nabla h({\bm{x}}^{t-1})$ § momentum
@@
For our purposes we can take ${\bm{d}}^{t}_{k}={\bm{g}}_{h}({\bm{y}}^{t}_{k-1},\xi^{t,k}_{h})+{\bm{m}}^{t}_{f-h}$ which should be a good approximation of $\nabla f({\bm{y}}^{t}_{k-1})$ and ${\bm{m}}^{t}_{f-h}$ is a momentum of fhf-h, we will consider two options: classical momentum or MVR (for momentum based variance reduction (MVR)).
Decentralized auxiliary information. More generally, we can assume having access to $N$ auxiliary functions $h_{i}({\bm{x}}):=\mathbb{E}_{\xi_{h_{i}}}\big[h_{i}({\bm{x}};\xi_{h_{i}})\big]$. While we can treat this case by taking $h=(1/N)\sum_{i=1}^{N}h_{i}$, we propose a more interesting solution that also works if the helpers $h_{i}$ are decentralized and cannot live in the same server. In this case, we can sample a set $S^{t}$ of helpers, each $h_{i}\in S^{t}$ will do the updates exactly as in Algorithm1, but this time ${\bm{x}}^{t}$ will be constructed using all $\{{\bm{y}}^{t}_{i,K},i\in S^{t}\}$. In our case, we propose to use the average ${\bm{x}}^{t}=(1/S)\sum_{i\in S^{t}}{\bm{y}}^{t}_{i,K}$ .
@@ 3 Algorithms and Results
Naive approach. This approach is exactly as suggested by its name, naive; it simply ignores the part fhf-h or, in other words, sets ${\bm{m}}^{t}_{f-h}=0$. The main idea is to use gradients (or gradient estimates) of $f$ whenever they are available and use gradients of $h$ when gradients of $f$ are not available. In our case, we alternate between one step using a gradient of $f$ and (K1)(K-1)-steps using gradients of $h$ without any correction. We will show that this approach suffers heavily from the bias between the gradients of $f$ and $h$. It is worth noting that in federated learning, this approach corresponds to Federated averaging (MacMahan2017).We remind the reader that we can take ${\bm{d}}^{t}_{k}={\bm{g}}_{h}({\bm{y}}^{t}_{k-1},\xi^{t,k}_{h})+{\bm{m}}^{t}_{f-h}$.
@@
where ${\bm{m}}_{f-h}^{t}$ is a momentum of fhf-h. We consider two methods for defining this momentum:
@@
AuxMOM simply uses the classical momentum, whereas AuxMVR uses the momentum-based variance reduction technique introduced in (MVR).
@@
In Assumption 3.2, we assume that we directly have access to unbiased gradient estimates of fhf-h; this does not restrict in any way our work since ${\bm{g}}_{f}-{\bm{g}}_{h}$ is such an estimate; however, this last estimate has a variance of $\sigma_{f-h}^{2}=\sigma_{f}^{2}+\sigma_{h}^{2}$, in general, it is possible to have a correlated estimate such that $\sigma_{f-h}^{2}<\sigma_{f}^{2}+\sigma_{h}^{2}$. If the batches $\xi_{f}$ and $\xi_{h}$ are drawn such that ${\bm{g}}_{f}$ and ${\bm{g}}_{h}$ are positively correlated, then it is even possible to have $\sigma_{f-h}^{2}<\sigma_{f}^{2}$.
@@
Discussion of the assumptions. Assumptions 3.1 and 3.2 are very common assumptions in the optimization literature. Under these two assumptions, it is well-known that SGD (SGd51; SGD52) has an optimal convergence rate. Assumption 3.3 was used extensively in Federated Learning (MiMe; karimireddy2019scaffold) and was considered for personalization (LinSpeedup).
Relaxing the Hessian similarity. GeneralizedSmoothness propose a generalized smoothness assumption where the norm of the Hessian can grow with the norm of the gradient. We believe it possible to extend our theory to accommodate such an assumption. In the same spirit, we can let $\|\nabla^{2}f({\bm{x}})-\nabla^{2}h({\bm{x}})\|_{2}$ grow with $\|\nabla f({\bm{x}})\|_{2}$. More important than this is finding similarity measures that apply to some of the potential applications we cite in Section 5.
@@ 4.1 Naive approach
In this section, we show the convergence results using the naive approach that uses a gradient of $f$ followed by K1K-1 gradients of $h$. For the analysis of this case (and only of this case), we need to make an assumption on the gradient bias between $f$ and $h$.
@@
There exists $f$ and $h$ satisfying assumptions 3.1,3.2,3.3,4.1 with $\delta=0$, $\sigma_{f}=\sigma_{h}=0$ such that $\frac{1}{KT}\sum_{t=1}^{T}\|\nabla f({\bm{x}}^{t-1})\|_{2}^{2}=\Omega(\zeta^{2})$. Using the biased SGD analysis in (BiasedSGD), it is easy to prove an upper bound, but we only need a lower bound for our purposes. Theorem 4.1 shows that this naive approach cannot guarantee convergence to less than $\zeta^{2}$ (up to some constant).
Note. Theorem 4.1 is very loose for Federated Averaging as, in this case, the function $f$ is directly tied to the helper entities (it is the average). Nevertheless, it’s worth noting that heterogeneity and client drift, which are akin to gradient bias, are recognized as factors that can constrain the performance of FedAVG. Consequently, there have been efforts to mitigate these issues through approaches such as those discussed in (karimireddy2019scaffold; MiMe).